# Dislocation correlations in GaN epitaxial films revealed by EBSD and XRD

### v1

@@ Abstract
Correlations between dislocations in crystals reduce the elastic energy via screening of the strain by the surrounding dislocations. We study the correlations of threading dislocations in GaN epitaxial films with dislocation densities of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2 by X-ray diffraction (XRD) in reciprocal space and by high-resolution electron backscatter diffraction (EBSD) in real space, where the strain is derived from a cross-correlation analysis of the Kikuchi patterns. The measured XRD curves and EBSD strain and rotation maps are compared with Monte Carlo simulations within one and the same model for the dislocation distributions. The screening of the dislocation strains is provided by creating pairs of dislocations with opposite Burgers vectors, with the mean distance between dislocations in a pair equal to the screening distance. The pairs overlap and cannot be distinguished as separate dipoles. The EBSD-measured autocorrelation functions of the strain and rotation components follow the expected logarithmic law for distances smaller than the screening distances and become zero for larger distances, which is confirmed by the Monte Carlo simulations. Screening distances of 2 µm and 0.3 µm are obtained for the samples with low and high dislocation densities, respectively. The dislocation strain is thus screened by only 4 neighboring dislocations. High-resolution EBSD allows for a more precise determination of the screening distances than from fits of the XRD curves. In addition, an anisotropic resolution of the EBSD measurements is observed and quantified.
@@ 1 Introduction
It is well established that the elastic energy of a single straight dislocation per unit length of the dislocation line $E\propto(\mu b^{2}/4\pi)\ln(R/r_{c})$ diverges as the lateral crystal size $R$ is increased to infinity. Here $\mu$ is the shear modulus, $b$ is the length of the Burgers vector, and $r_{c}$ is the dislocation core radius ($r_{c}\simeq 2.6b$) [1]. Hence, the elastic energy density of a crystal containing a finite density of random uncorrelated dislocations tends to infinity with increasing crystal size. Wilkens proposed [2] that the dislocations in a crystal are correlated, thereby screening the long range strain field from each other. He proposed a model of a “restrictedly random dislocation distribution”, where the crystal is subdivided into cells, with each cell containing the same number of dislocations. The numbers of dislocations with opposite Burgers vectors in a cell are equal, so that the total Burgers vector of a cell is zero and the total strain from the dislocations in a cell decays with the distance faster than the strain of an individual dislocation (in Wilkens’ model, the strain is assumed to be restricted within the cell). Then, the cell size $R$ replaces the crystal size in the expression above for the elastic energy, and the energy density becomes independent of the crystal size.
@@
This model was proposed in the framework of X-ray diffraction (XRD) line profile analysis of polycrystals, since the same divergence affects the widths of the diffraction lines. The strain probability density distribution calculated for screw dislocations in the Wilkens model [3] is used in computer programs for XRD line profile analysis [4, 5]. The fits of the experimental curves involve two parameters of the dislocation distribution, the dislocation density $\varrho$ and the cell size $R$. Although the initial formulation of the model implies that the cell size $R$ is large compared to the mean distance between dislocations $\varrho^{-1/2}$ (so that the number of the dislocations in a cell is large), in many cases these distances are comparable. A Monte Carlo simulation of the XRD profiles literally implementing the Wilkens model leads to artifacts when a cell contains one or only a few dislocation pairs [6]. An alternative model of the dislocation strain screening is to consider random pairs of dislocations with opposite Burgers vectors, with $R$ being the mean distance between dislocations in a pair [7]. When the mean separation in a pair $R$ exceeds the mean distance between all dislocations $\varrho^{-1/2}$, the pairs overlap and cannot be distinguished as individual dipoles.
In many XRD studies, the dislocation density $\varrho$ is of primary interest, while the screening distance $R$ serves as an auxiliary fit parameter required for an accurate determination of the dislocation density. However, the elastic energy stored in a crystal with dislocations is also of interest and requires an accurate determination of $R$ [8]. In the XRD studies, the diffracted intensity is collected from the whole sample and the screening distance $R$ is accessed only as a transition from a Gaussian shaped central part of the diffraction line to its $\propto q^{-3}$ tails. Recently, the scanning XRD microscopy technique has been developed to obtain maps of spatial distributions of strain and rotation tensors [9]. Such a study can be performed on a synchrotron and provides the strain averaged over the film thickness, due to the large penetration depth of the X-rays. This averaging does not affect the maps for the case of threading dislocations in (In,Ga)N films studied in Ref. [9], but for complicated three-dimensional dislocation distributions it may blur the maps up to a complete loss of information.
@@
High-resolution electron backscatter diffraction (EBSD) provides maps of spatial distributions of the strain and rotation tensor components near the crystal surface obtained from shifts of the Kikuchi bands [10, 11, 12]. The strain sensitivity of the method reached $10^{-4}$ with the invention of the a cross-correlation based analysis of EBSD patterns [13, 14]. The strain and rotation maps and the strain probability density distributions have been reported in several studies [15, 16, 17]. The maps obtained in the present study are similar to those reported in the previous studies. However, the primary aim of our work is to obtain and analyze the spatial correlations of the strains, which was not addressed in these previous studies. We show that the strain-strain correlations calculated from the maps clearly reveal the dislocation strain screening by surrounding dislocations and provide the screening distances $R$.
@@
The analysis of dislocations in GaN and other group III nitrides is of primary interest for optoelectronic applications of these materials and, at the same time, can serve as a touchstone for various methods of evaluating dislocation arrays. GaN epitaxial films contain threading dislocations oriented parallel to each other and perpendicular to the film surface. Such well-defined dislocation arrays allow a detailed modeling of both XRD curves and EBSD maps and their quantitative comparison. A respective comparison of different methods for dislocation density determination in metals [18] is much more complicated since it requires an account of several slip systems, grain boundaries, etc.
@@
We carry out both XRD and EBSD measurements on the same samples and also perform Monte Carlo simulations of the XRD line profiles and EBSD strain maps, using one and the same model of dislocation distributions for both experimental techniques and the same computing program. A comparison of the measured and Monte-Carlo-modeled strain maps in EBSD provides us with two further findings in addition to the determination of the screening distance for the dislocation strain. First, we find that the resolution of EBSD is highly anisotropic, with significantly worse resolution in the direction of the inclination of the incident electron beam. A quantitative estimate of the resolution is obtained by comparing the strain-strain correlations in two orthogonal directions. Secondly, we find that EBSD underestimates the strain for large dislocation densities in the sample grown by molecular beam epitaxy. This finding is considered to be the result of the dynamical diffraction of electrons in a crystal exhibiting a large strain gradient in the region where the diffracted beam is formed.
@@ 2.1 Samples
We study three GaN(0001) samples with different densities of threading dislocations (see Fig. 1). In all samples, the dislocations are straight lines running from the substrate to the film surface, along the surface normal. Hence, threading dislocations provide well defined arrays in single crystal samples suitable for both XRD and EBSD studies. The investigated samples were characterized in our previous works [19, 20, 21].
As a reference, we use a 350 µm-thick free-standing GaN (0001) film grown by hydride vapor phase epitaxy. We refer to it as sample 0. The dislocation density in this film is estimated from cathodoluminescence images to be as low as $6\times 10^{5}$ cm-2. Since this sample produced an unexpectedly broad strain distribution in the EBSD measurements described in Sec. 3.2 below, it was rinsed with isopropanol and ethanol and then cleaned with oxygen plasma to remove any contaminants and produce a uniformly oxidized surface layer. Etching of the surface with heated KOH removed this oxidized layer and left a smooth surface with a root-mean-square roughness of 0.31 nm. However, this cleaning step had very little effect on the EBSD maps.
@@
Additionally, a reference EBSD measurement is obtained on a Si(001) wafer.
@@ 2.3 Electron backscatter diffraction
Scanning electron microscopy (SEM) and CL imaging, as well as the EBSD measurements were carried out in a Zeiss Ultra 55 scanning electron microscope. For the EBSD measurements, the microscope was operated at an acceleration voltage of 15 kV with a beam current of 6 nA. To record the Kikuchi patterns, the stage with the mounted sample was tilted by 70 with respect to the incident beam direction towards an EDAX Hikari Super EBSD detector.
The patterns were recorded with the highest available resolution of $470\times 470$ pixels and background-corrected using a reference pattern. EBSD maps were recorded over either $2\times 2$ µm2 or $10\times 10$ µm2 areas with step sizes of 20 nm and 50 nm, respectively, at an exposure time of 100 ms. Measurements with these settings took approximately 20 min for the smaller maps and 80 min for the bigger maps to record. Sample drift was estimated by comparing SEM images recorded before and after the measurement and were smaller than 18% of the map dimensions, indicating negligible drifts per recorded pixel but a potential warping of the spatial resolution.
The cross-correlation analysis of the Kikuchi patterns [^1] Early studies distinguish Kikuchi patterns formed by electron diffraction in transmission electron microscopy, and pseudo-Kikuchi patterns due to diffraction of backscattered electrons in SEM [31]. Nowadays, the term Kikuchi pattern is used for both of them. [32] was carried out with the software CrossCourt 4.5.3.6 by BLG Vantage. Examples of Kikuchi patterns from all four samples are presented in Fig. S1 in the Supplementary Material (SM). 20 regions of interest on the Kikuchi patterns were referenced for the calculations of the cross-correlation function. From these calculations, maps of the strain and rotation tensor components with a sub-pixel resolution were obtained using the elastic moduli of GaN [33]. A remapping of the resulting maps had a minor influence on the strain distribution and was not applied for the shown data.
@@ 2.4 Monte Carlo simulations
Monte Carlo simulations of both XRD curves and EBSD maps are performed in the framework of one and the same model of dislocation arrangements and dislocation strains using the computing programs that were developed earlier for XRD [19, 34] and adapted here for EBSD. The displacement fields of a-type edge dislocations with Burgers vectors $1/3\left\langle 11\bar{2}0\right\rangle$, c-type screw dislocations with Burgers vectors $\left\langle 0001\right\rangle$, and a+c-type mixed dislocations with Burgers vectors $1/3\left\langle 11\bar{2}3\right\rangle$ are calculated. The screening of the dislocation strains by surrounding dislocations is provided by generating dislocations in pairs with opposite Burgers vectors. The positions of the pairs are random and uncorrelated. The distances between dislocations in the pairs are assumed to follow a lognormal distribution with a mean value of $R$ (hereafter referred to as the screening distance) and a standard deviation of $R/2$. The XRD intensity $I(q)$ is calculated as a probability density of the distortion component $q=\hat{\mathbf{K}}{}^{\mathrm{out}}\cdot\nabla(\mathbf{Q}\cdot\mathbf{U})$, where $\mathbf{U}(\mathbf{r})$ is the total displacement due to all dislocations, $\mathbf{Q}$ is the diffraction vector, and $\hat{\mathbf{K}}{}^{\mathrm{out}}$ is the unit vector in the direction of the diffracted beam [19, 34]. It is worth noting that the direction $\hat{\mathbf{K}}{}^{\mathrm{out}}$ emerges in the diffraction from a single crystal as a consequence of the intensity integration performed by a widely open detector over the plane in reciprocal space perpendicular to $\mathbf{K}^{\mathrm{out}}$. In powder diffraction, the respective intensity integration from randomly oriented grains is performed over the plane perpendicular to the diffraction vector. In this case, $\hat{\mathbf{K}}{}^{\mathrm{out}}$ is replaced with the unit vector in the direction of $\mathbf{Q}$ [30]. The EBSD maps are simulated by calculating all components of the strain tensor $\varepsilon_{ij}=(\partial U_{i}/\partial x_{j}+\partial U_{j}/\partial x_{i})/2$ and the rotation tensor $\omega_{ij}=(\partial U_{i}/\partial x_{j}-\partial U_{j}/\partial x_{i})/2$, using the same displacement vector $\mathbf{U}(\mathbf{r})$. For the XRD simulations, it is sufficient to use the displacement field of a straight dislocation in an infinite crystal, as the X-ray intensity is collected over the entire film thickness. The displacement for a dislocation along the $\left\langle 0001\right\rangle$ direction in a hexagonal crystal is employed. It coincides with the isotropic solution with the Poisson ratio taken to be $\nu_{h}=c_{12}/(c_{11}+c_{12})$ ([35], Sec. 2.5). Here $c_{ij}$ are the elastic moduli. Using their values for GaN [33], we obtain $\nu_{h}=0.27$. For the EBSD simulations, the elastic strain relaxation at the free surface is taken into account. We have compared the isotropic solution ([36], Sec. 6.3) with a more comprehensive anisotropic solution for a dislocation along the sixfold axis ([35], Sec. 2.5). The difference is negligible compared to the statistical errors in the Monte Carlo simulations presented below.

### v2

@@ Abstract
Correlations between dislocations in crystals reduce the elastic energy via screening of the strain by the surrounding dislocations. We study the correlations of threading dislocations in GaN epitaxial films with dislocation densities of $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2 by X-ray diffraction (XRD) in reciprocal space and by high-resolution electron backscatter diffraction (HR-EBSD) in real space, where the strain is derived from a cross-correlation analysis of the Kikuchi patterns. The measured XRD curves and HR-EBSD strain and rotation maps are compared with Monte Carlo simulations within one and the same model for the dislocation distributions. The screening of the dislocation strains is modeled by creating pairs of dislocations with opposite Burgers vectors, with the mean distance between dislocations in a pair equal to the screening distance. The pairs overlap and cannot be distinguished as separate dipoles. The HR-EBSD-measured autocorrelation functions of the strain and rotation components follow the expected logarithmic law for distances smaller than the screening distances and become zero for larger distances, which is confirmed by the Monte Carlo simulations. The kink in the plot of the autocorrelation function allows a robust and accurate determination of the screening distance without making any simulation or fit. Screening distances of 2 µm and 0.3 µm are obtained for the samples with low and high dislocation densities, respectively. The dislocation strain is thus screened by only 4 neighboring dislocations. In addition, an anisotropic resolution of the HR-EBSD measurements is observed and quantified.
@@ 1 Introduction
It is well established that the elastic energy of a single straight dislocation per unit length of the dislocation line $E\propto(\mu b^{2}/4\pi)\ln(R/r_{c})$ diverges as the lateral crystal size $R$ is increased to infinity. Here $\mu$ is the shear modulus, $b$ is the length of the Burgers vector, and $r_{c}$ is the dislocation core radius ($r_{c}\simeq 2.6b$) [1]. Hence, the elastic energy density of a crystal containing a finite density of random uncorrelated dislocations tends to infinity as the crystal size increases. The energy remains finite when dislocations are correlated, such that the long-range strain field of a dislocation is screened by surrounding dislocations. Then, $R$ is the screening distance, rather than the crystal size.
@@
Both dislocation density and the screening distance affect the intensity distributions in X-ray diffraction (XRD). The XRD line profile analysis [2, 3] is an established method in powder diffraction for determining the dislocation density as well as the screening distance, and hence the elastic energy stored in a crystal, and changes to these quantities during plastic deformation. The calculation of the diffracted intensity requires a model for the dislocation arrangement. Wilkens proposed [4, 5] a “restrictedly random dislocation distribution”, where the crystal is subdivided into cells, with each cell containing the same number of dislocations. The numbers of dislocations with opposite Burgers vectors in a cell are equal so that the total Burgers vector of a cell is zero, resulting in the total strain from the dislocations in a cell decaying with the distance faster than the strain of an individual dislocation (in Wilkens’ model, the strain is assumed to be restricted within the cell). The cell size $R$ then plays the role of the screening distance for dislocation strains. The strain probability density distribution calculated for screw dislocations in the Wilkens model [5] is used in computer programs for XRD line profile analysis [6, 7].
@@
The XRD intensity distribution can be calculated for any given dislocation distribution model by means of Monte Carlo integration [8, 9]. A Monte Carlo simulation of the XRD profiles that implements the Wilkens model literally leads to artifacts when a cell contains only one or a few dislocation pairs [10]. An alternative model for dislocation strain screening is to consider random pairs of dislocations with opposite Burgers vectors, with $R$ being the mean distance between dislocations in a pair [11]. When $R$ exceeds the mean distance between all dislocations $\varrho^{-1/2}$, the pairs overlap and cannot be distinguished as individual dipoles. Below, we only deal with such overlapping pairs and do not consider the case of dislocation dipoles with the dipole widths small compared to the separation between dipoles.
In laboratory XRD studies of dislocations, the diffracted intensity is collected from the entire sample and integrated over the diffraction vector orientation in powder diffraction or the diffracted wave direction in a double-crystal diffraction setup for single crystals. The XRD line profile analysis is based on fitting of the experimental curves involving two parameters of the dislocation distribution, the dislocation density $\varrho$ and the screening distance $R$. In many XRD studies, the dislocation density $\varrho$ is of primary interest, while the screening distance $R$ is an auxiliary fit parameter required for accurately determining the dislocation density. However, determination of the elastic energy stored in a crystal requires an accurate determination of $R$ [12]. The screening distance $R$ is accessed as a transition from the Gaussian central part of the diffraction line to its tails, which show an intensity decay that is proportional to $q^{-3}$. These tails originate from diffraction in highly disturbed regions close to dislocation lines and their intensity is independent of dislocation correlations.
The screening distance $R$ can be accessed directly if spatial strain maps are measured. $R$ can then be obtained from the strain-strain correlation function [13]. In that work, all components of the strain and rotation tensors have been measured on the example of threading dislocations in (In,Ga)N films by the scanning XRD microscopy technique [14], using synchrotron X-rays. Mapping in three non-coplanar reflections was performed to obtain all nine components of the strain and rotation tensors. Since the dislocation strain depends on the distance from the dislocation line $r$ by the universal law $\varepsilon\propto r^{-1}$, the correlation function decreases proportionally to the logarithm of the distance for distances smaller than $R$, and becomes zero at larger distances. Plotting the correlation function against the logarithm of the distance yields a hook-shaped curve. The kink in this curve provides the screening distance without the need for any fitting or assumptions about the dislocation correlations.
A variety of lattice-sensitive microscopy techniques for mapping strain and rotation in crystals is nowadays available [15]. A number of synchrotron-based XRD techniques, such as X-ray Laue microdiffraction [16, 17], dark-field X-ray microscopy [18], scanning XRD microscopy [14], Bragg coherent diffraction imaging [19], or Bragg X-ray ptychography [20], provide access to the full strain tensor. The correlations of strain fields have been studied to characterize dislocation distributions by scanning XRD microscopy [13] and dark-field X-ray microscopy [21].
@@
While these XRD techniques are restricted to synchrotron radiation sources, high-resolution electron back scattering diffraction (HR-EBSD) is a widely available laboratory technique that provides spatial maps of all components of the strain and rotation tensors near the crystal surface. The strains and rotations are obtained from shifts of the Kikuchi bands [22, 23, 24]. The strain sensitivity of the method was enhanced to $10^{-4}$ with the invention of the a cross-correlation-based analysis of EBSD patterns [25, 26]. Strain and rotation maps, as well as the strain probability density distributions, have been reported in several studies [27, 28, 29]. The maps obtained in the present study are similar to those reported in the previous studies. However, the primary aim of our work is to obtain and analyze the spatial correlations of the strains, a subject not addressed in these previous studies. We demonstrate that the strain-strain correlations calculated from the maps clearly reveal the dislocation strain screening by surrounding dislocations and provide the screening distances $R$.
@@
The analysis of dislocations in GaN and other group III nitrides is of primary interest for optoelectronic applications of these materials and, at the same time, can serve as a touchstone for various methods of evaluating dislocation arrays. GaN epitaxial films contain threading dislocations oriented parallel to each other and perpendicular to the film surface. These well-defined dislocation arrays enable detailed modeling of both XRD curves and HR-EBSD maps and their quantitative comparison. A respective comparison of different methods for dislocation density determination in metals [30] is much more complicated since it requires an account of several slip systems, grain boundaries, etc.
@@
We carry out both laboratory XRD and HR-EBSD measurements on the same samples and also perform Monte Carlo simulations of the XRD line profiles and HR-EBSD strain maps, using one and the same model of dislocation distributions for both experimental techniques and the same computing program. Comparing the measured and Monte-Carlo-modeled strain maps in HR-EBSD provides us with two further findings in addition to the determination of the screening distance for the dislocation strain. First, we find that the resolution of HR-EBSD is highly anisotropic, with significantly worse resolution in the direction of the inclination of the incident electron beam. We obtain a quantitative estimate of the resolution by comparing the strain-strain correlations in two orthogonal directions. Secondly, we find that the experimental HR-EBSD maps show, as uncorrelated random noise, shear strain components that should be zero at a strain-free surface. The widths of their distributions can be used as an estimate of the accuracy of the HR-EBSD measurements.
@@ 2.1 Samples
We study three GaN(0001) samples with different densities of threading dislocations (see Fig. 1). In all samples, the dislocations are straight lines running from the substrate to the film surface, along the surface normal. Hence, threading dislocations provide well defined arrays in single crystal samples suitable for both XRD and HR-EBSD studies. The investigated samples were characterized in our previous works [31, 32, 33].
As a reference, we use a 350 µm-thick free-standing GaN (0001) film grown by hydride vapor phase epitaxy. We refer to it as sample 0. The dislocation density in this film is estimated from cathodoluminescence images to be as low as $6\times 10^{5}$ cm-2. Since this sample produced an unexpectedly broad strain distribution in the HR-EBSD measurements described in Sec. 3.2 below, it was rinsed with isopropanol and ethanol and then cleaned with oxygen plasma to remove any contaminants and produce a uniformly oxidized surface layer. Etching of the surface with heated KOH removed this oxidized layer and left a smooth surface with a root-mean-square roughness of 0.31 nm. However, this cleaning step had very little effect on the HR-EBSD maps.
@@
Additionally, a reference HR-EBSD measurement is obtained on a Si(001) wafer.
@@ 2.3 Electron backscatter diffraction
Scanning electron microscopy (SEM) and CL imaging, as well as the HR-EBSD measurements were carried out in a Zeiss Ultra 55 scanning electron microscope. For the HR-EBSD measurements, the microscope was operated at an acceleration voltage of 15 kV with a beam current of 6 nA. To record the Kikuchi patterns, the stage with the mounted sample was tilted by 70 with respect to the incident beam direction towards an EDAX Hikari Super EBSD detector.
The patterns were recorded with the highest available resolution of $470\times 470$ pixels and background-corrected using a reference pattern. HR-EBSD maps were recorded over either $2\times 2$ µm2 or $10\times 10$ µm2 areas with step sizes of 20 nm and 50 nm, respectively, at an exposure time of 100 ms. Measurements with these settings took approximately 20 min for the smaller maps and 80 min for the bigger maps to record. Sample drift was estimated by comparing SEM images recorded before and after the measurement and were smaller than 18%. Its effect is quantified and discussed at the end of Sec. 3.3, as the discussion is based on the results of Monte Carlo simulations. We find that the effect of drift can be neglected.
Cross-correlation analysis of the Kikuchi patterns [^1] Early studies distinguish Kikuchi patterns formed by electron diffraction in transmission electron microscopy, and pseudo-Kikuchi patterns due to diffraction of backscattered electrons in SEM [43]. Nowadays, the term Kikuchi pattern is used for both of them. [44] was carried out with the software CrossCourt 4.5.3.6 from BLG Vantage. Examples of Kikuchi patterns from all four samples are presented in Fig. SM1 in the Supplementary Material (SM). The software defaults were chosen for 20 regions of interest (squared subsets of the Kikuchi pattern). A reference pattern was chosen as the pixel with the lowest kernel average misorientation in the initial indexation of the EBSD map. The Kikuchi pattern of each region was compared to the same area of a reference pattern to identify tiny shifts and rotations of the pattern. The cross-correlation method enables these changes to be determined with a sub-pixel resolution. From changes in different parts of the Kikuchi pattern, the strain and rotation tensor components are obtained using the elastic moduli of GaN [45]. Applying this analysis to every point of an EBSD map enables the changes in strain and crystal rotation to be mapped.
@@ 2.4 Monte Carlo simulations
Monte Carlo simulations of both XRD curves and HR-EBSD maps are performed in the framework of one and the same model of dislocation arrangements and dislocation strains using the computing programs that were developed earlier for XRD [31, 46] and adapted here for HR-EBSD. We include in the simulations all types of perfect threading dislocations that are present in wurtzite GaN epitaxial layers [47, 48, 49, 50]. The displacement fields of a-type edge dislocations with Burgers vectors $1/3\left\langle 11\bar{2}0\right\rangle$, c-type screw dislocations with Burgers vectors $\left\langle 0001\right\rangle$, and a+c-type mixed dislocations with Burgers vectors $1/3\left\langle 11\bar{2}3\right\rangle$ are calculated. Crystallographically equivalent Burgers vectors are taken with equal probabilities.
@@
The screening of the dislocation strains by surrounding dislocations is provided by generating dislocations in pairs with opposite Burgers vectors. The positions of the pairs are random and uncorrelated. The distances between dislocations in the pairs are taken from a lognormal distribution with a mean value of $R$ (hereafter referred to as the screening distance) and a standard deviation of $R/2$. The simulations are insensitive to the standard deviation: a unimodal distribution (all dislocation pairs have the same distance between dislocations) gives the same result. The XRD intensity $I(q)$ is calculated as a probability density of the distortion component $q=\hat{\mathbf{K}}{}^{\mathrm{out}}\cdot\nabla(\mathbf{Q}\cdot\mathbf{U})$, where $\mathbf{U}(\mathbf{r})$ is the total displacement due to all dislocations, $\mathbf{Q}$ is the diffraction vector, and $\hat{\mathbf{K}}{}^{\mathrm{out}}$ is the unit vector in the direction of the diffracted beam [31, 46]. It is worth noting that the direction $\hat{\mathbf{K}}{}^{\mathrm{out}}$ emerges in the diffraction from a single crystal as a consequence of the intensity integration performed by a widely open detector over the plane in reciprocal space perpendicular to $\mathbf{K}^{\mathrm{out}}$. In powder diffraction, the respective intensity integration from randomly oriented grains is performed over the plane perpendicular to the diffraction vector. In this case, $\hat{\mathbf{K}}{}^{\mathrm{out}}$ is replaced with the unit vector in the direction of $\mathbf{Q}$ [42].
The HR-EBSD maps are simulated by calculating all components of the strain tensor $\varepsilon_{ij}=(\partial U_{i}/\partial x_{j}+\partial U_{j}/\partial x_{i})/2$ and the rotation tensor $\omega_{ij}=(\partial U_{i}/\partial x_{j}-\partial U_{j}/\partial x_{i})/2$, using the same displacement vector $\mathbf{U}(\mathbf{r})$. For the XRD simulations, it is sufficient to use the displacement field of a straight dislocation in an infinite crystal, as the X-ray intensity is collected over the entire film thickness. The displacement for a dislocation along the $\left\langle 0001\right\rangle$ direction in a hexagonal crystal is employed. It coincides with the isotropic solution with the Poisson ratio taken to be $\nu_{h}=c_{12}/(c_{11}+c_{12})$ ([51], Sec. 2.5). Here $c_{ij}$ are the elastic moduli. Using their values for GaN [45], we obtain $\nu_{h}=0.27$. For the HR-EBSD simulations, the elastic strain relaxation at the free surface is taken into account. We have compared the isotropic solution ([52], Sec. 6.3) with a more comprehensive anisotropic solution for a dislocation along the sixfold axis ([51], Sec. 2.5). The difference is negligible compared to the statistical errors in the Monte Carlo simulations presented below.